# SurgLaVi: Large-Scale Hierarchical Dataset for Surgical Vision-Language Representation Learning

### v1

@@ 3.4 Stage 4: Contextual Caption Enrichment
In the final stage, we enhance filtered captions by incorporating temporal context and metadata to improve semantic richness and coherence. For each clip–caption pair , we apply with a designed prompt (Fig. 10 in SM) to instruct the model to generate an enriched caption conditioned on: (ii) a fixed window of preceding captions at the same granularity level, (iii) the video title, and (iv) the surgical procedure type.
@@ 3.5 SurgLaVi Dataset Overview
Video Source Collection. To construct SurgLaVi, we aggregated surgical videos from public and private sources and processed them through our automated pipeline. For the open-source subset, SurgLaVi-, we leveraged prior large-scale web scrapes of surgical YouTube videos provided by LEMON [18] and GenSurgery [19], incorporating 3,110 and 4,172 videos, respectively, for a total of 6,812 unique videos. To further expand the scale, diversity, and domain coverage of the full dataset, SurgLaVi additionally includes 2,945 videos sourced from private collections of surgical academic lectures, resulting in a total of 9,758 videos.
@@
Quantitative summary. After processing the video collection through our pipeline, the final SurgLaVi dataset consists of 5,317 videos totaling 843 hours of surgical footage, sampled at varying frame rates from 10 to 60 fps (mean: 29 fps). The dataset contains 239,776 clip–caption pairs organized across three hierarchical granularities: phase, step, and task level (see Fig. 4 for the distribution across levels). After filtering, the phase-level clips, the base level of our hierarchical structure, comprise 664.9 hours of video with 70M frames. SurgLaVi- represents the open-source subset of SurgLaVi, containing 2,464 videos with 465 hours of surgical footage, sampled at 15 to 30 fps (mean: 28.75 fps), and 112,757 clip–caption pairs following the same hierarchical structure. The filtered phase-level clips in SurgLaVi- total 398 hours of video with 41M frames.
@@
Hierarchical structure illustration. To qualitatively demonstrate the hierarchical organization of our dataset, Figure 3 presents a portion of a representative video segmented into phase, step, and task levels as defined in Stage 2. Moving from phase to task level results in segments of progressively shorter duration and higher temporal resolution, while the corresponding captions evolve from long, context-rich descriptions to concise, action-focused statements. In the illustration, each clip is represented with the same number of sampled frames, highlighting how the effective sampling rate increases across levels: phase clips provide broader coverage with lower temporal resolution, while task clips capture fine-grained details with higher temporal resolution. This design explicitly encodes multiple temporal scales, enabling models trained on SurgLaVi to perform multi-scale temporal reasoning over surgical activities.
@@
Hierarchical level distribution. Figure 5 illustrates the distribution of clip durations (left) and the number of clips per video (right) across hierarchical levels. In both SurgLaVi and SurgLaVi-, phase-level clips span longer intervals with relatively few segments per video, step-level clips provide intermediate coverage, and task-level clips are shorter but occur more frequently. Despite differences in video sources, the distributions of clip durations and clip counts per level remain consistent between the two datasets. Complementing these temporal statistics, Figure 4 presents the overall distribution of clip–caption pairs across levels. As granularity increases from phase to task, the number of clips rises accordingly, a trend observed in both the full SurgLaVi dataset and its open-source subset, SurgLaVi-.
Comparison with prior datasets. Table 1 contrasts SurgLaVi and SurgLaVi with existing surgical video–caption datasets. Compared to SVL and GenSurg+, our datasets offer three key advantages: (i) Scale: our full and open-source versions contain substantially more videos and clip–caption pairs; (ii) Hierarchical temporal segmentation: prior datasets use clips of uniform duration, whereas ours provide clips of varying lengths to capture procedural context at multiple granularities (phase, step, task); (iii) Accessibility: our dataset includes an openly available subset with captions, while all previous datasets are either entirely private (SVL) or have private captions (GenSurg+).
@@
SurgLaVi | 5,317 | P: 35.4 | P: 86.6 | ✗ | ||||||
S: 64.2 | S: 45.1 | |||||||||
T: 140.2 | T: 18.2 | |||||||||
F: 239.8 | F: 28.6 | |||||||||
SurgLaVi | 2,464 | P: 15.7 | P: 65.4 | ✓ | ||||||
S: 30.0 | S: 36.1 | |||||||||
T: 66.9 | T: 15.2 | |||||||||
F: 112.6 | F: 35.6 |
Dataset diversity. Figure 6 depicts the distribution of videos in SurgLaVi and SurgLaVi across three taxonomy levels: surgical specialty, subject, and specific procedure type. The full dataset covers 10 specialties, 39 subjects, and 217 procedure types, while the open-source subset spans 9 specialties, 34 subjects, and 170 procedure types. For comparison, the GenSurg+ dataset includes only 28 procedure types (although the specific types differ from those in our taxonomy). We used an internal taxonomy designed by a team of expert surgeons (see Fig. 3 from SM) to define a hierarchical classification scheme from specialty down to specific procedure type and assigned labels in a weakly supervised manner by providing GPT-4o with the taxonomy and each video’s transcription summary and prompting it to classify the video at all three levels (Fig. 11 in SM).
@@ 4 SurgCLIP
We establish a base model for our dataset using a dual-encoder video–language model trained with a symmetric contrastive objective. The video encoder is a ViT-B/16 backbone, used in the form of a TimeSFormer model [20] with divided space–time attention, which decouples spatial and temporal attention blocks. This design offers flexibility during inference, allowing the same model to operate on different temporal inputs, ranging from a single frame to multi-frame clips, without re-training. The text encoder is a BERT-base model [21] that processes the caption tokens. Both encoders project their modality-specific features into a shared embedding space via learned projection heads.
Given the hierarchical nature of our dataset, we randomly sample clip–caption pairs from different levels (phase, step, task) within each training batch, as shown in Figure 7. This strategy ensures that the InfoNCE objective is computed over a mixed-level pool of samples, enabling cross-level alignment and encouraging the model to learn representations transferable across temporal granularities. Specifically, for a batch , where is a video clip, its text caption, and its granularity level, we encode them as: , where and denote the video and text encoders, respectively. The video-to-text and text-to-video InfoNCE losses across all levels are:
@@
Model | Phase recognition | Step | Action | Tool | ||||||||||||||||||||||
Cholec80 | AutoLaparo | StrasBypass70 | BernBypass70 | GraSP | GraSP | SARRARP50 | CholeT50 | Cholec80 | GraSP | |||||||||||||||||
Acc | F1 | Acc | F1 | Acc | F1 | Acc | F1 | Acc | F1 | Acc | F1 | Acc | F1 | mAP | mAP | mAP | ||||||||||
CLIP | 27.81 | 8.42 | 8.02 | 4.79 | 18.52 | 3.51 | 20.73 | 4.17 | 9.93 | 2.06 | 3.85 | 0.83 | 28.80 | 6.28 | 2.50 | 18.44 | 36.00 | |||||||||
MedSigLIP | 40.94 | 18.74 | 26.54 | 10.17 | 20.15 | 5.54 | 30.73 | 6.14 | 12.88 | 4.84 | 11.53 | 1.86 | 5.49 | 3.17 | 2.98 | 17.92 | 37.18 | |||||||||
SurgVLP | 49.02 | 32.46 | 10.02 | 7.19 | 26.08 | 19.18 | 30.46 | 17.48 | 11.79 | 6.07 | 6.02 | 3.22 | 17.91 | 7.03 | 3.04 | 31.19 | 36.93 | |||||||||
HecVL | 47.50 | 23.89 | 38.72 | 25.81 | 28.59 | 23.91 | 32.09 | 18.58 | 10.99 | 6.46 | 1.96 | 0.56 | 4.67 | 3.03 | 3.63 | 25.11 | 36.75 | |||||||||
PeskaVLP | 51.43 | 39.39 | 34.94 | 28.41 | 29.42 | 19.64 | 31.48 | 17.47 | 13.85 | 9.86 | 3.65 | 2.47 | 5.18 | 2.68 | 4.49 | 38.88 | 41.05 | |||||||||
SurgCLIP(β) | 57.98 | 39.42 | 55.72 | 45.95 | 31.24 | 26.05 | 18.30 | 15.06 | 34.77 | 27.98 | 14.15 | 11.14 | 13.94 | 7.62 | 4.17 | 36.77 | 43.06 | |||||||||
SurgCLIP | 61.29 | 50.53 | 69.14 | 56.37 | 32.37 | 30.78 | 23.90 | 19.68 | 41.49 | 34.94 | 26.28 | 16.53 | 17.42 | 7.76 | 5.28 | 40.80 | 45.97 |
By computing similarities across all pairs in the mixed batch, we implicitly optimize the model for cross-level retrieval, making it sensitive to fine-grained and coarse-grained procedural semantics. To further support this objective, we apply a dynamic temporal sampling strategy that adapts the temporal resolution of clips to their hierarchical level. While the number of frames per clip remains fixed, the sampling stride varies with the clip’s duration. Task-level clips, short segments focused on fine-grained surgical actions, are sampled at higher frame rates, providing high temporal resolution to capture subtle instrument motions and action-centric captions. In contrast, phase-level clips cover longer temporal intervals and are sampled at lower frame rates, producing temporally coarser representations aligned with broader procedural descriptions. This combination of mixed-level batching and dynamic temporal sampling exposes the model to multi-scale temporal structure, enabling it to capture fine-grained manipulations and broader procedural context within a unified pre-training framework.
@@ 5.1 Zero-Shot classification
Experimental setting and benchmarks. We evaluate generalization and transferability in a zero-shot setting across laparoscopic and robotic modalities, multiple procedure types, and tasks at different granularities. For phase recognition, we use Cholec80 [22], cholecystectomy), AutoLaparo [23], hysterectomy), StrasByPass and BernByPass [24], gastric bypass), and GraSP [25], robotic prostatectomy). We evaluate step recognition on GraSP [25], action recognition on SARRARP50 [26], robotic prostatectomy), and triplet (instrument–verb–target) recognition on CholecT50 [27]. We also assess tool presence recognition on Cholec80 and GraSP. We compare against vision–language models trained on natural images (CLIP, [1], medical images (MedSigLIP, [28], and surgical video (SurgVLP, [6]; HecVL, [9]; PeskaVLP, [8]). For fairness, we use each method’s original prompts when available.
@@
Inference setting. All models follow a consistent zero-shot protocol. The frozen vision backbone encodes a 16-frame window centered on each evaluation frame to model temporal dynamics in surgical videos. Then, we mean pool features into a single embedding and compare via cosine similarity to all candidate class embeddings. The top-similarity class is predicted for phase, step, and action recognition, and we report video-wise Accuracy and F1-score following Twinanda et al. [22]. For tool presence and CholecT50 triplets, we treat similarity scores as class probabilities and report mAP. We provide inference prompts specific for each dataset and task in the supplementary material.
@@
Results. Table 2 shows that our approach consistently outperforms prior work across nearly all datasets and tasks. Notably, SurgCLIP pre-trained on SurgLaVi-, a smaller subset, already surpasses baselines on most benchmarks, while pre-training on the full SurgLaVi results in further gains. This pattern indicates that our automated curation pipeline produces high-quality supervision from open-source videos, and that additional scale, diversity, and hierarchical structure provide incremental benefits and robust zero-shot transfer across procedures, modalities, and task granularities. The largest gains occur in high-level temporal tasks such as phase recognition, with improvements of up to +10, +28, and +25 F1 over the strongest prior methods on Cholec80, AutoLaparo, and GraSP, respectively. While action recognition remains particularly challenging, SurgCLIP matches or exceeds previous VLP approaches and delivers consistent gains in tool recognition, indicating that its learned representations extend beyond workflow understanding to object-centric tasks. A remaining limitation, however, is that current models, including SurgCLIP, primarily rely on global visual features during pre-training and inference, which limits their capacity to capture the fine spatial details and subtle instrument motions necessary for precise action recognition.
5.2 Linear Probing
Experimental setting and benchmarks. To evaluate the quality and transferability of our learned representations, we perform few- and full-shot linear probing by training a linear classifier on frozen features from SurgCLIP’s vision backbone. We report results on three benchmarks: Cholec80 and AutoLaparo for surgical phase recognition, and GraSP for phase and step recognition. Following our zero-shot setting, we compare SurgCLIP against CLIP, MedSigLIP, and PeskaVLP for linear probing. In addition, we include LEMON [18], a large self-supervised surgical foundation model, to explicitly assess the added value of language supervision during pre-training relative to purely visual objectives. All models are probed using mean-pooled features from 32-frame windows at inference, following our analysis on temporal window size impact (Fig. 8).
Few-shot and full-shot linear probing. We evaluate linear probing in 10%, 50%, and 100% training regimes. For the 10% setting, we generate five random subsets of the training data without replacement and report the mean and standard deviation across runs. For the 50% setting, we also generate five random subsets, where overlap across subsets may occur but each subset is sampled without replacement. The 100% regime uses the full training set in a single run. This protocol tests robustness under limited supervision while ensuring consistency across models.
Results. As shown in Figure 8, SurgCLIP outperforms all baselines across every regime trained on SurgLaVi and SurgLaVi-. We attribute this to three design choices: (1) a video backbone that explicitly models temporal dynamics, (2) language supervision that aligns features with procedural semantics, and (3) hierarchical pre-training with dynamic temporal sampling, which exposes the model to multi-scale surgical activity. Together, these generate semantically enriched representations that are discriminative even when only a linear layer is trained on top. This consistent advantage highlights several trends. First, transferring features from general-purpose models such as CLIP is challenging, as the large domain gap to surgery reduces performance despite massive pre-training scale. Second, within-domain pre-training provides clear benefits: PeskaVLP leverages surgical video–language data but is limited by dataset size and a lightweight backbone, while MedSigLIP benefits from larger-scale medical vision–language supervision and surpasses PeskaVLP despite not being specialized for surgery. Third, LEMON illustrates the strength of large-scale self-supervision on surgical frames, producing strong spatial features but without explicitly modeling temporal dynamics or procedural semantics. Finally, although full SurgLaVi pre-training achieves the strongest results, the competitiveness of SurgLaVi- highlights the effectiveness of our automated curation pipeline, which produces high-quality supervision even from a smaller data scale.

### v2

@@ 3.4 Stage 4: Contextual Caption Enrichment
In the final stage, we enhance filtered captions by incorporating temporal context and metadata to improve semantic richness and coherence. For each clip–caption pair , we apply with a designed prompt (Fig. 10 in Appendix) to instruct the model to generate an enriched caption conditioned on: (i) a fixed window of preceding captions at the same granularity level, (ii) the video title, and (iii) the surgical procedure type.
@@ 3.5 SurgLaVi Dataset Overview
Video Source Collection. To construct SurgLaVi-, we leveraged prior large-scale web scrapes of surgical YouTube videos provided by LEMON [che2025lemon] and GenSurgery [gensurg], incorporating 3,110 and 4,172 videos, respectively, for a total of 6,812 unique videos. These videos originate from surgeons and institutions that publicly share operative recordings for educational purposes. To further expand the scale, diversity, and domain coverage of the full dataset, SurgLaVi, we additionally include 2,945 videos sourced from a surgical training archive that is part of an institutionally approved educational repository and adheres to ethical standards for surgical training. These materials do not contain identifiable patient information, and their use for AI research has been formally reviewed and authorized. This brings the total video source size to 9,758 videos.
@@
Quantitative summary. After processing the video collection through our pipeline, the final SurgLaVi dataset consists of 5,317 videos totaling 843 hours of surgical footage, sampled at varying frame rates from 10 to 60 fps (mean: 29 fps). The dataset contains 239,776 clip–caption pairs organized across three hierarchical granularities: coarse, mid, and fine level (see Fig. 4 for the distribution across levels). After filtering, the coarse-level clips, the base level of our hierarchical structure, comprise 664.9 hours of video with 70M frames. SurgLaVi- represents the open-source subset of SurgLaVi, containing 2,464 videos with 465 hours of surgical footage, sampled at 15 to 30 fps (mean: 28.75 fps), and 112,757 clip–caption pairs following the same hierarchical structure. The filtered coarse-level clips in SurgLaVi- total 398 hours of video with 41M frames.
@@
Hierarchical structure illustration. To qualitatively demonstrate the hierarchical organization of our dataset, Figure 3 presents a portion of a representative video segmented into coarse, mid, and fine levels. Moving from coarse to fine level results in segments of progressively shorter duration and higher temporal resolution, while the corresponding captions evolve from long, context-rich descriptions to concise, action-focused statements. In the illustration, each clip is represented with the same number of sampled frames, highlighting how the effective sampling rate increases across levels: coarse clips provide broader coverage with lower temporal resolution, while fine clips capture fine-grained details with higher temporal resolution. This design explicitly encodes multiple temporal scales, enabling models trained on SurgLaVi to perform multi-scale temporal reasoning over surgical activities.
@@
Hierarchical level distribution. Figure 5 illustrates the distribution of clip durations (left) and the number of clips per video (right) across hierarchical levels. In both SurgLaVi and SurgLaVi-, coarse-level clips span longer intervals with relatively few segments per video, mid-level clips provide intermediate coverage, and fine-level clips are shorter but occur more frequently. Despite differences in video sources, the distributions of clip durations and clip counts per level remain consistent between the two datasets. Complementing these temporal statistics, Figure 4 presents the overall distribution of clip–caption pairs across levels. As granularity increases from coarse to fine, the number of clips rises accordingly, a trend observed in both the full SurgLaVi dataset and its open-source subset, SurgLaVi-.
@@
| SurgLaVi | 5,317 | C: 35.4 | C: 86.6 | ✗ | ||||||
| M: 64.2 | M: 45.1 | |||||||||
| F: 140.2 | F: 18.2 | |||||||||
| A: 239.8 | A: 28.6 | |||||||||
| SurgLaVi | 2,464 | C: 15.7 | C: 65.4 | ✓ | ||||||
| M: 30.0 | M: 36.1 | |||||||||
| F: 66.9 | F: 15.2 | |||||||||
| A: 112.6 | A: 35.6 |
Comparison with prior datasets. Table 1 contrasts SurgLaVi and SurgLaVi with existing surgical video–caption datasets. Compared to SVL and GenSurg+, our datasets offer three key advantages: (i) Scale: our full and open-source versions contain substantially more videos and clip–caption pairs; (ii) Hierarchical temporal segmentation: prior datasets use clips of uniform duration, whereas ours provide clips of varying lengths to capture procedural context at multiple granularities (coarse, mid, fine); (iii) Accessibility: our dataset includes an openly available subset with captions, while all previous datasets are either entirely private (SVL) or have private captions (GenSurg+).
Dataset diversity. Figure 6 depicts the distribution of videos in SurgLaVi and SurgLaVi across three taxonomy levels: surgical specialty, subject, and specific procedure type. The full dataset covers 10 specialties, 39 subjects, and 217 procedure types, while the open-source subset spans 9 specialties, 34 subjects, and 170 procedure types. For comparison, the GenSurg+ dataset includes only 28 procedure types (although the specific types differ from those in our taxonomy). We used an internal taxonomy designed by a team of expert surgeons (see Fig. 6 from Appendix) to define a hierarchical classification scheme from specialty down to specific procedure type and assigned labels in a weakly supervised manner by providing GPT-4o with the taxonomy and each video’s transcription summary and prompting it to classify the video at all three levels (Fig. 11 in Appendix).
@@ 4 SurgCLIP
We establish a base model for our dataset using a dual-encoder video–language model trained with a symmetric contrastive objective. The video encoder is a ViT-B/16 backbone, used in the form of a TimeSFormer model [bertasius2021space] with divided space–time attention, which decouples spatial and temporal attention blocks. This design offers flexibility during inference, allowing the same model to operate on different temporal inputs, ranging from a single frame to multi-frame clips, without re-training. The text encoder is a BERT-base model [devlin2019bert] that processes the caption tokens. Both encoders project their modality-specific features into a shared embedding space via learned projection heads.
Given the hierarchical nature of our dataset, we randomly sample clip–caption pairs from all hierarchical levels (coarse, mid, fine) within each training batch, as shown in Figure 7. This strategy ensures that the InfoNCE objective is computed over a mixed-level pool of samples, enabling cross-level alignment and encouraging the model to learn representations transferable across temporal granularities. Specifically, for a batch , where is a video clip, its text caption, and its granularity level, we encode them as: , where and denote the video and text encoders, respectively. The video-to-text and text-to-video InfoNCE losses across all levels are:
@@
By computing similarities across all pairs in the mixed batch, we implicitly optimize the model for cross-level retrieval, making it sensitive to fine-grained and coarse-grained procedural semantics. To further support this objective, we apply a dynamic temporal sampling strategy that adapts the temporal resolution of clips to their hierarchical level. While the number of frames per clip remains fixed, the sampling stride varies with the clip’s duration. Fine-level clips, short segments focused on fine-grained surgical actions, are sampled at higher frame rates, providing high temporal resolution to capture subtle instrument motions and action-centric captions. In contrast, coarse-level clips cover longer temporal intervals and are sampled at lower frame rates, producing temporally broader representations aligned with procedural descriptions. This combination of mixed-level batching and dynamic temporal sampling exposes the model to multi-scale temporal structure, enabling it to capture fine-grained manipulations and broader procedural context within a unified pre-training framework.
| Recognition Task | Dataset | Metric | CLIP | MedSigLIP | SurgVLP | HecVL | PeskaVLP | SurgCLIP(β) | SurgCLIP |
| Phase | Cholec80 | Acc | 27.8110.4 | 40.94±1.14 | 49.02±10.6 | 47.50±11.0 | 51.43±11.01 | 57.98±5.25 | 61.29±10.7 |
| F1 | 8.42±2.81 | 18.74±4.42 | 32.46±6.43 | 23.89±5.53 | 39.39±7.43 | 39.42±5.25 | 50.53±8.83 | ||
| AutoLaparo | Acc | 8.02±3.43 | 26.54±5.39 | 10.02±5.23 | 38.72±7.01 | 34.94±9.06 | 55.72±7.65 | 69.14±5.98 | |
| F1 | 4.79±1.77 | 10.17±3.24 | 7.19±4.15 | 25.81±4.91 | 28.41±7.79 | 45.95±8.10 | 56.37±4.54 | ||
| StrasBypass70 | Acc | 18.52±5.99 | 20.15±3.98 | 26.08±6.30 | 28.59±8.52 | 29.42±6.92 | 31.24±6.88 | 32.37±6.30 | |
| F1 | 3.51±1.12 | 5.54±1.15 | 19.18±4.30 | 23.91±6.60 | 19.64±3.97 | 26.05±4.39 | 30.78±5.37 | ||
| Heichole | Acc | 19.21±5.27 | 45.36±10.7 | 45.70±10.7 | 29.91±12.9 | 53.33±8.47 | 56.95±17.59 | 63.84±15.4 | |
| F1 | 6.17±2.43 | 17.63±3.24 | 27.04±3.92 | 22.76±7.21 | 36.46±4.43 | 44.00±7.68 | 55.15±8.9 | ||
| BernBypass70 | Acc | 20.73±9.14 | 30.73±8.13 | 30.46±6.55 | 32.09±8.41 | 31.48±8.60 | 18.30±6.68 | 23.90±6.58 | |
| F1 | 4.17±1.23 | 6.14±1.34 | 17.48±2.93 | 18.58±5.00 | 17.47±5.25 | 15.06±5.54 | 19.68±5.78 | ||
| GraSP | Acc | 9.93±4.79 | 12.88±7.20 | 11.79±2.89 | 10.99±2.99 | 13.85±3.45 | 34.77±8.15 | 41.49±7.87 | |
| F1 | 2.06±0.57 | 4.84±2.30 | 6.07±1.71 | 6.46±2.87 | 9.86±3.14 | 27.98±6.53 | 34.94±7.61 | ||
| Step | GraSP | Acc | 3.85±2.72 | 11.53±6.63 | 6.02±1.11 | 1.96±1.54 | 3.65±1.85 | 14.15±3.92 | 26.28±7.29 |
| F1 | 0.83±0.61 | 1.86±1.01 | 3.22±0.25 | 0.56±0.38 | 2.47±0.80 | 11.14±4.16 | 16.53±5.44 | ||
| Action | SARRARP50 | Acc | 28.8±3.85 | 5.49±3.24 | 17.91±2.86 | 4.67±3.09 | 5.18±3.95 | 13.94±8.18 | 17.42±7.08 |
| F1 | 6.28±1.14 | 3.17±1.75 | 7.03±1.99 | 3.03±2.43 | 2.68±1.87 | 7.62±2.91 | 7.76±2.71 | ||
| Triplet | CholeT50 | mAP | 2.50 | 2.98 | 3.04 | 3.63 | 4.49 | 4.17 | 5.28 |
| Tool | Cholec80 | mAP | 18.44 | 17.92 | 31.19 | 25.11 | 38.88 | 36.77 | 40.80 |
| Heichole | 21.18 | 20.58 | 22.46 | 19.36 | 32.82 | 31.47 | 36.79 | ||
| GraSP | 36.00 | 37.18 | 36.93 | 36.75 | 41.05 | 43.06 | 45.97 |
For detailed implementation details about training schedules, hyperparameters and computational resources used to train SurgCLIP, refer to Sec. 1.2 in Appendix.
@@ 5.1 Zero-Shot Classification
Experimental setting and benchmarks. We evaluate generalization and transferability in a zero-shot setting across laparoscopic and robotic modalities, multiple procedure types, and a wide range of temporal granularities. This includes recognition levels that can match our pre-training as well as benchmarks that operate at even more granular level, such as action, triplet, and tool recognition. Evaluating these additional levels enables us to measure how well the model handles tasks whose temporal granularity exceeds supervision during pretraining. For phase recognition, we cover five laparoscopic benchmarks: Cholec80 [endonet] and HeiChole [wagner2023comparative] for cholecystectomy, AutoLaparo [wang2022autolaparo] for hysterectomy, and StrasByPass and BernByPass [multibypass] for gastric bypass—as well as the robotic GraSP benchmark [ayobi2025pixel] for robotic prostatectomy. For step recognition, we use GraSP [ayobi2025pixel]. Action recognition is evaluated on SARRARP50 [sarrarp50], which focuses on robotic prostatectomy. Triplet recognition (instrument–verb–target) is assessed using CholecT50 [nwoye2022rendezvous], and tool presence recognition is evaluated on Cholec80 [endonet], GraSP [ayobi2025pixel], and Heichole [wagner2023comparative].
@@
Inference setting. All models follow a consistent zero-shot protocol. The frozen vision backbone encodes a 16-frame window centered on each evaluation frame to model temporal dynamics in surgical videos. Then, we mean pool features into a single embedding and compare via cosine similarity to all candidate class embeddings. For phase, step, and action recognition, the predicted label corresponds to the class with the highest similarity score. We evaluate performance using video-wise Accuracy and video-wise macro-averaged F1 as proposed in [funke2023metrics] and report standard deviation across videos. For the video-wise macro-averaged F1, we first compute the F1-score for each class within a procedure and then average these scores uniformly across classes. Next, we average the resulting per-procedure F1-scores across all procedures to avoid bias toward longer videos and to better capture variability between cases. For simplicity, we refer to these metrics as F1 and Accuracy throughout the remainder of this work. For tool presence and CholecT50 triplets, we treat similarity scores as class probabilities and report mAP. We provide inference prompts specific for each dataset and task in the Appendix.
@@
Results. Table 2 shows that our approach consistently outperforms prior work across nearly all datasets and tasks. Notably, SurgCLIP pre-trained on SurgLaVi-, a smaller subset, already surpasses baselines on most benchmarks, while pre-training on the full SurgLaVi results in further gains. This pattern indicates that our automated curation pipeline produces high-quality supervision from open-source videos, and that additional scale, diversity, and hierarchical structure provide incremental benefits and robust zero-shot transfer across procedures, modalities, and task granularities. The largest gains occur in high-level temporal tasks such as phase recognition, with improvements of up to +10, +28, and +25 F1 over the strongest prior methods on Cholec80, AutoLaparo, and GraSP, respectively. While action recognition remains particularly challenging, SurgCLIP matches or exceeds previous VLP approaches and delivers consistent gains in tool recognition, indicating that its learned representations extend beyond workflow understanding to object-centric tasks. A remaining limitation, however, is that models primarily rely on global visual features during pre-training and inference, which limits their capacity to capture the fine spatial details and subtle instrument motions necessary for precise action recognition.
5.2 Linear Probing and Context Optimization
Experimental setting and benchmarks. To evaluate the quality and transferability of the learned representations, we conduct Linear Probing and Context Optimization (CoOp) [zhou2022learning]. On the one hand, linear probing assesses the discriminative capabilities of vision-only representations by training a single linear classifier on frozen visual embeddings from SurgCLIP’s video backbone. On the other hand, CoOp probes multimodal alignment by learning a set of continuous context tokens that condition the text encoder, enabling us to assess how effectively each model’s visual and textual representations interact when guided by task-specific prompts. Together, these two evaluations provide a comprehensive view of representation quality in unimodal and multimodal settings.
We conduct these studies on three downstream benchmarks: Cholec80 and AutoLaparo for surgical phase recognition, and GraSP for phase and step recognition. Consistent with our zero-shot analysis, we compare SurgCLIP against CLIP, MedSigLIP, and PeskaVLP. Additionally, we include LEMON [che2025lemon], a surgical foundation model, to compare our approach against in-domain, large-scale, self-supervised, purely visual pre-training strategies. All models are evaluated using frozen features and identical inference settings to ensure strict comparability.
5.2.1 Linear Probing
We consider two complementary few-shot paradigms that reflect different low-effort labeling scenarios:
(1) Video Few- and Full-Shot Linear Probing. This paradigm follows prior surgical VLP work [peskavlp, surgvlp, hecvl]. Here, N-shot denotes the percentage of labeled procedure videos (i.e., full video procedures) sampled from the training split. Once the video of a patient is selected, all frames from that procedure are used to train the linear classifier. We evaluate 10%, 50%, and 100% regimes, where the latter corresponds to full-shot linear probing. This setting reflects a practical annotation scenario in which only a limited number of patient video recordings can be labeled, and annotations are provided at phase/step transition times, which allows all frames within each segment to automatically inherit the corresponding label.
(2) CLS Few-Shot Linear Probing: Inspired by few-shot evaluation protocols in general computer vision [zhang2022tip, zhou2022learning], we introduce a second paradigm in which few-shot supervision is defined at the frame level rather than the procedure level. Unlike the video-based regime, where selecting a video provides all frames from that procedure, this setting imposes a stricter constraint by limiting the actual number of annotated samples, not the number of procedures. This also increases training diversity, as sampled frames may originate from multiple surgeries, reflecting a scenario where many patient videos are available but only a small number of frames per class can be annotated. In this setting, we sample 1, 2, 4, 8, and 16 labeled samples per class from the full training split.
Setting. All models are trained and evaluated using mean-pooled embeddings from a 32-frame temporal window, following our analysis of temporal context (Fig. 9). For each labeled frame, we construct this window by gathering 32 unlabeled neighboring frames around it. This preserves the supervision at a single annotated frame while allowing the model to leverage local temporal structure. For all few-shot studies, we generate five random subsets of the training data without replacement and report the mean and standard deviation across runs, ensuring that identical subsets are used across all models for strict comparability. More implementation details available in Appendix, Sec. 1.3.1.
5.2.2 Context Optimization (CoOp)
For multimodal evaluation, we apply Context Optimization (CoOp) [zhou2022learning] using a single shared learnable context across all classes. Instead of hand-crafted prompts, each class is represented only by its class name, and the learnable context tokens are prepended to this name to form the final text embedding. The vision and text encoders remain frozen, ensuring that CoOp measures the adaptability of the prompt space rather than the encoders themselves.
We follow CoOp’s few-shot training protocol, aligned with our CLS Few-Shot Linear Probing setup. Specifically, we sample 1, 4, 8, and 16 labeled frames per class from the training split and optimize only the context tokens. As with linear probing, we construct five random subsets for each shot level and report the mean and standard deviation across runs. Additional implementation details in Appendix, Sec. 1.3.1.
Results. As shown in Figure 9, SurgCLIP consistently outperforms all baselines across CoOp, CLS Few-Shot, and Video Few-/Full-Shot settings for both SurgLaVi and SurgLaVi-. This advantage holds across nearly all datasets and supervision regimes, with the only exception occurring in GraSP step recognition under CoOp. We attribute this performance to three design choices: (1) a video backbone that explicitly models temporal dynamics, (2) language supervision that aligns features with procedural semantics, and (3) hierarchical pre-training with dynamic temporal sampling, which exposes the model to multi-scale surgical activity. Together, these components produce semantically enriched representations that remain highly discriminative even when only a linear layer is trained on top.
The results also reveal several consistent trends. First, CoOp improves upon zero-shot inference when more than two labeled samples are available, demonstrating the benefit of learnable contextual prompts over hand-crafted text embeddings; however, it generally remains less effective than linear probing on the visual features. Secondly, using only 16 labeled frames per class often outperforms training on all frames from a single labeled video, and in several cases (AutoLaparo–50%, GraSP–50%, Cholec80–10%) approaches the performance of labeling 4–5 full procedures. This highlights the practical value of sparse, diverse annotations over dense single-procedure labeling. Third, the relative performance of competing models varies across datasets: on Cholec80, PeskaVLP is the strongest non-SurgCLIP baseline, while on AutoLaparo and GraSP, LEMON generally performs best among non-SurgCLIP models. Interestingly, MedSigLIP, despite being trained on broad medical data rather than surgical data, often matches or exceeds PeskaVLP, underscoring the benefit of large-scale medical image–text supervision. Conversely, CLIP is consistently the lowest-performing model, reflecting the substantial domain gap between natural images and surgical videos despite its massive pre-training scale.